# SF-Loc: A Visual Mapping and Geo-Localization System based on Sparse Visual Structure Frames

### v1

@@ Abstract
For high-level geo-spatial applications and intelligent robotics, accurate global pose information is of crucial importance. Map-aided localization is an important and universal approach to overcome the limitations of global navigation satellite system (GNSS) in challenging environments. However, current solutions face challenges in terms of mapping flexibility, storage burden and re-localization performance. In this work, we present SF-Loc, a lightweight visual mapping and map-aided localization system, whose core idea is the map representation based on sparse frames with dense (though downsampled) depth, termed as visual structure frames. In the mapping phase, multi-sensor dense bundle adjustment (MS-DBA) is applied to construct geo-referenced visual structure frames. The local co-visbility is checked to keep the map sparsity and achieve incremental mapping. In the localization phase, coarse-to-fine vision-based localization is performed, in which multi-frame information and the map distribution are fully integrated. To be specific, the concept of spatially smoothed similarity (SSS) is proposed to overcome the place ambiguity, and pairwise frame matching is applied for efficient and robust pose estimation. Experimental results on both public and self-made datasets verify the effectiveness of the system. In complex urban road scenarios, the map size is down to 3 MB per kilometer and stable decimeter-level re-localization can be achieved. The code will be made open-source soon (https://github.com/GREAT-WHU/SF-Loc).
@@
Dense bundle adjustment, multi-sensor fusion, visual localization and mapping
@@ I Introduction
Global localization is an important issue in robotics applications and autonomous driving [1, 2]. For outdoor scenarios, global navigation satellite system (GNSS) is a commonly used technology for achieving global drift-free localization[3]. However, precise GNSS relies on high-precision services, while low-cost GNSS can hardly provide stable and accurate localization solutions in complex environments[4]. Map matching is another method for achieving global localization. By associating observations of the surrounding environment from external sensors widely equipped on intelligent devices (such as cameras) with a priori maps, different levels of global localization can be achieved, depending on the map representation[5, 6].
To achieve the practicality of map-aided localization, both the mapping and the re-localization phases should be looked into. As to the mapping phase, the efficiency of map construction, the accuracy of the map elements and the storage burden of map data are some of the most important factors. Considering the need of incremental mapping and crowd-sourced map updating, the flexibility of the map representation is also crucial. Current map forms, like vector map[7], 3D point cloud[8, 9], occupancy grid[10] and neural representations[11, 12], still cannot meet the above requirements perfectly, especially when considering vision-based schemes.
@@
In this work, we propose SF-Loc, a lightweight visual mapping and map-aided localization system based on sparse visual structure frames. The visual structure frame is defined as a geo-tagged image frame with compressed RGB map and full-view depth map, which itself supports standalone and flexible re-localization. The word “structure” comes from two aspects, as the frame stores the environmental structure information (in the depth form), and the map frames are sparsely structured with special consideration of the geometric distribution for efficient association and low redundancy. Based on the concept, SF-Loc provides a pipeline for the auto-generation of visual structure frames through multi-sensor dense bundle adjustment (MS-DBA), as well as an implementation for coarse-to-fine map-based localization. The contributions of this work are listed as follows:
@@
3) Real-world, cross-season experiments are conducted to validate different phases of the system, and the data sequences are made public.
4) The code is made open-source to benefit the community.
@@ II-A VSLAM with multi-sensor fusion
Visual SLAM, which has been extensively studied in the past decade, is considered an effective and low-cost solution for relative pose estimation and mapping. Considering the form of visual information, point features[8, 13], line features[14] and direct photometric errors[15] have been exploited. By integrating IMU[16, 17, 18], GNSS[19, 3], and other sensors[20], the practicality of VSLAM can be improved in terms of robustness, continuity, accuracy and geo-referencing functionality. On this basis, multi-session and multi-agent SLAM systems are developed[21, 22]. Learning-based VSLAM frameworks are also investigated in recent implementations, covering neural representations[23, 24], data-driven association[25, 26] and end-to-end pipelines[27, 28].
For map reuse of multi-sensor integrated VSLAM, obtaining globally smoothed pose estimation in the geographic coordinate system is crucial, as it impacts the absolute accuracy of the map and the practicality of user-end localization.
@@ II-B Visual place recognition
Visual place recognition (VPR) is a long-existing problem, which plays a crucial role in loop closure, multi-session mapping, re-localization, etc. Early approaches use handcrafted models (like bag-of-words[29] and VLAD[30]) to aggregate local feature descriptors, thus to indicate the scene similarity. To overcome the long-term change challenge, sequential information is utilized[31].
In the last decade, DNN-based methods, like NetVLAD[32], have greatly improved the practicality and accuracy of VPR. Following this trend, TransVPR[33] uses the transformer to incorporate multi-level attention mechanisms and improve the VPR performance. CosPlaces[34] transforms the place recognition problem from contrastive learning to a classification problem, which enables efficient training on 10-million-level large-scale datasets and is further improved with better viewpoint robustness[35]. AnyLoc[36] utilizes the large visual foundation model to achieve ubiquitous VPR for different scenarios. It is noted that, most existing methods don’t consider comprehensively the spatial distribution of database images and user-end relative poses, which could serve as powerful information to improve the stability of VPR.
@@ II-C Fine-grained visual re-localization
For high-accuracy visual re-localization (e.g. decimeter- or centimeter-level), it is generally needed to perform feature- or instance-level association between user-end perception and the map elements. Many of these methods follow a coarse-to-fine pipeline. In ORB-SLAM3[8] and VINS-Mono[16], re-localization is performed through bag-of-words-based place recognition and pointwise matching between new feature observations and the reconstructed feature point cloud. Such pipeline is further improved in HLoc[9], where the power of deep learning is fully utilized in both mapping and localization phases. Besides, literature[38] proposes a map-free visual re-localization method by exploiting monocular depth prediction.
Along with these methods, other approaches try to exploit higher-level information for map matching. Instance-level object modeling is integrated into classic VSLAM systems to overcome viewpoint variation and achieve long-term consistency[39, 40, 41]. In recent works, bird’s-eye view features are utilized to associate with priori vector maps to achieve accurate pose estimation, both for outdoor[42] and indoor scenarios[43]. The tradeoff of semantic-based re-localization is the loss of low-level, sometimes subtle but distinguishable, visual cues, which might lead to higher ambiguity.
@@ III System Overview
In the mapping phase, multi-sensor information is used to generate the point-free map frames with recovered dense depth. Global co-visibility graph optimization is performed periodically in a low frequency to achieve global optimal pose estimation. On this basis, the local co-visibility is checked to keep the map sparsity and to achieve incremental mapping. We also employ a DNN-based VPR model to compute the global descriptors of map frames.
In the localization phase, deep visual place recognition and fine-grained feature matching are utilized to achieve coarse-to-fine localization. During this process, multi-frame information and the map frame distribution are fully considered to improve the recall rate and the accuracy.
One principle behind the proposed map representation is to keep the visual information raw and non-encoded. The first advantage of doing so is that it ensures users have flexibility when using the map, without being constrained by any intermediate forms of information. The second advantage is that, the compression of a raw image is effective by applying mature techniques like JPEG, without losing important geometric or appearance information.
@@ IV-A Multi-Sensor Dense Bundle Adjustment
The concept of dense bundle adjustment (DBA) is proposed in[26], which is inherited in this work. Specifically, sequential images are fed into a CNN with gated recurrent units (GRUs) to compute dense optical flows among co-visible image pairs. To explain this, consider the rigid flow produced by dense projection of one image pair (source frame $i$ projected to target frame $j$)
@@
where $\Pi_{c}$ is the camera projection model, $\mathbf{T}_{ij}$ is the relative pose, $\mathbf{u}_{i}$ denotes the $n=(H/8)\times(W/8)$ grid-like 2D points in the image frame, $\mathbf{\lambda}_{i}$ is the inverse depth map.
The optical flow module takes the image features and the rigid flow as the input and outputs the corrected optical flow (in the residual form) together with the weight $\mathbf{w}_{ij}$, which leads to the following error equation
@@
The residual-form dense optical flows are treated as re-projection errors, which are further used to construct a bundle adjustment problem to simultaneously estimate the poses and dense depths of the images. Furthermore, the estimated poses and depths are recurrently fed back into the optical flow module, thus to iteratively refine all the geometric information. This process is end-to-end trainable, which shows good generalization in real-world scenes with synthetic training[26].
In our early work[45], we demonstrate the possibility of tightly integrating DBA with multi-sensor information. To be specific, the Hessian form of (2) is derived as and stacked for a window of poses around frame $i$
@@
where $\bm{\xi}_{i,1,\cdots,N}$ is the stacked camera pose Lie algebras.
@@
$\mathbf{x}_{k}=\left[\begin{matrix}\mathbf{T}^{w}_{b_{k}}&\mathbf{v}^{w}_{b_{k}}&\mathbf{b}_{k}\end{matrix}\right],k\in\mathcal{K}$ (5)
$\mathbf{T}^{w}_{b_{k}}=\left[\begin{matrix}\mathbf{R}^{w}_{b_{k}}&\mathbf{t}^{w}_{b_{k}}\\ 0&1\end{matrix}\right]\in SE(3),\ \ \mathbf{b}_{k}=\left[\begin{matrix}\mathbf{b}_{a,k}&\mathbf{b}_{g,k}\end{matrix}\right]$ (6)
@@
where $\mathbf{H}_{c}$ and $\mathbf{v}_{c}$ are the Hessian matrix and vector of DBA with all the depth states eliminated, $\mathbf{x}_{c}$ is the related camera poses. To associate the camera poses with the IMU-centered pose states, the IMU-camera extriniscs are used to transform (7), as detailed in[45].
@@
$\min\mathbf{E}_{m}(\mathbf{x})+\mathbf{E}_{c}(\mathbf{x}_{c})+\sum_{k\in\mathcal{K}}\left\|\mathbf{r}_{b}(\mathbf{x}_{k},\mathbf{x}_{k+1})\right\|+\sum_{k\in\mathcal{G}}\left\|\mathbf{r}_{g}(\mathbf{x}_{k})\right\|$ (10)
where $\mathbf{E}_{m}(\mathbf{x}_{k})$ is the marginalization term in the quadric cost function form.
@@
and the estimated poses and depths are fed back into (1) and (2) to refine the optical flows and adjust the weights.
The optical flow module will be recurrently called for several times until mature, which simultaneously adjust the pose estimation, the dense data association and the depths, trying to achieve consistency of the above information.
An issue that should be considered is the obtainment of global optimal (or approximately) pose estimation. This is important both for monocular VIO (due to the dynamic-dependent observability) and GNSS-integrated navigation (due to intermittent GNSS availability). In our system, we maintain a global factor graph, which collects the marginalized factors from the sliding-window estimator. The objective of the global factor graph is written as follows
$\min\!\!\sum_{k\in\mathcal{K}_{\text{all}}}\!\!\mathbf{E}_{c,k}(\mathbf{x}_{c,k})\!+\!\!\sum_{k\in\mathcal{K}_{\text{all}}}\!\!\left\|\mathbf{r}_{b}(\mathbf{x}_{k},\mathbf{x}_{k+1})\right\|\!\!+\!\!\sum_{k\in\mathcal{G}_{\text{all}}}\!\!\left\|\mathbf{r}_{g}(\mathbf{x}_{k})\right\|$ (12)
@@ IV-B Map Construction and Maintenance
After the global pose optimization, we extract the frame pose from the global graph, pack the RGB image (with JPEG compression), the low-resolution depth map and the global descriptor produced by some VPR model together to construct a visual structure frame, then insert it into the map. To make the map lightweight, it is important to keep the sparsity of the visual structure frames without sacrificing too much of the effectiveness of user-end re-localization. Besides, considering the application scenario of multi-session mapping, the system should support incremental update.
@@ V-A Multi-Frame Place Recognition
We use lightweight CNN-based methods (e.g., [32, 36, 49, 34, 35]) to compute global descriptors and to evaluate the similarity between the query image and map frames. On this basis, we develop a model-driven technique to integrate sequential queries and map frame distribution to improve the robustness and stability of the coarse localization.
@@
To avoid the ambiguity of single-shot VPR, we propose an metric called spatially smoothed similarity (SSS) that combines the information of appearance similarity, map distribution and user-end relative poses. To be specific, we set particles around every landmark frames in the map, with every particle to be a 2D pose $(x,y,\theta)$ that represents a candidate of the query pose. With the help of the user-end odometry, we compute the virtual trajectory of every particle, following
$\left(\mathbf{T}^{i}\right)_{4\times 4\times W}=\mathbf{T}^{w}_{p_{i}}\left\{\mathbf{T}^{c_{k}}_{c_{k+1-j}}\right\}_{j=1,2,\cdots,W},{i\in\mathcal{P}}$ (15)
where $\mathcal{P}$ is the particle candidate set, $W$ is the window size used for multi-frame VPR.
@@
$\left({\text{idx}}^{i}\right)_{W\times 1}=\text{db}.\text{nearest}(\mathbf{T}^{i}),\ i\in\mathcal{P}$ (16)
$\left(\bm{\sigma}^{i}\right)_{W\times 1}=\bm{\sigma}\left[\text{idx}^{i}\right],\ i\in\mathcal{P}$ (17)
where db is the map database organized by a spatial KD-Tree, $\bm{\sigma}$ is the $W\times M$ similarity matrix which stores the L2 similarities between the $W$ query frames and all $M$ frames in the map. The similarity computation is performed every time a new query frame comes and stored in the memory.
Then, we compute the SSS by averaging the L2 distances
$\sigma^{i}_{\text{SSS}}=||\bm{\sigma}^{i}||_{2},\ i\in\mathcal{P}$ (18)
@@
The rationality of SSS is based on the assumption that the historical images of a correctly recalled query should also exhibit high similarity to their spatially adjacent map frames. The geographical distribution of map frames (which is obtained in the mapping phase), along with user-end pose information, is implicitly incorporated to the metric in nearest neighbor search. This technique can be easily adopted to common VPR models without extra layers or finetuning, transforming single-frame VPR similarities to multi-frame ones.
@@
Figure 5: Illustration of the multi-frame place recognition. Particles that indicate the query pose are set around map frames. The user-end trajectory is used for nearest searching to associate with the map frames, then the SSS is evaluated, which combines the information of multiple image pairs.

### v2

@@ Abstract
For high-level geo-spatial applications and intelligent robotics, accurate global pose information is of crucial importance. Map-aided localization is a universal approach to overcome the limitations of global navigation satellite system (GNSS) in challenging environments. However, current solutions face challenges in terms of mapping flexibility, storage burden and re-localization performance. In this work, we present SF-Loc, a lightweight visual mapping and map-aided localization system, whose core idea is the map representation based on sparse frames with dense but compact depth, termed as visual structure frames. In the mapping phase, multi-sensor dense bundle adjustment (MS-DBA) is applied to construct geo-referenced visual structure frames. The local co-visbility is checked to keep the map sparsity and achieve incremental mapping. In the localization phase, coarse-to-fine vision-based localization is performed, in which multi-frame information and the map distribution are fully integrated. To be specific, the concept of spatially smoothed similarity (SSS) is proposed to overcome the place ambiguity, and pairwise frame matching is applied for efficient and robust pose estimation. Experimental results on the cross-season dataset verify the effectiveness of the system. In complex urban road scenarios, the map size is down to 3 MB per kilometer and stable decimeter-level re-localization can be achieved.
@@
Dense bundle adjustment, multi-sensor fusion, visual localization and mapping, place recognition
@@ I Introduction
Global localization is a critical problem in robotics and autonomous driving applications [1, 2]. For outdoor scenarios, global navigation satellite system (GNSS) is a commonly used technology for achieving global drift-free localization[3]. However, precise GNSS generally relies on high-precision services and products, while low-cost GNSS can hardly provide stable and accurate localization solutions in complex environments[4]. Map matching is another method for achieving global localization. By associating observations of the surrounding environment from external sensors widely equipped on intelligent devices (such as cameras) with a priori maps, different levels of global localization can be achieved, depending on the map representation[5, 6].
To achieve the practicality of map-aided localization, both the mapping and the re-localization phases should be carefully addressed. As to the mapping phase, the efficiency of map construction, the accuracy of the map elements and the storage burden of map data are some of the key factors. Considering the need of incremental mapping and crowd-sourced map updating, the flexibility of the map representation is also crucial. Current map forms, like vector map[7], 3D point cloud[8, 9], occupancy grid[10] and neural representations[11, 12], still cannot meet the above requirements perfectly, especially when considering vision-based schemes.
@@
In this work, we propose SF-Loc, a lightweight visual mapping and map-aided localization system based on sparse visual structure frames. The visual structure frame is defined as a geo-tagged image frame with compressed RGB map and full-view depth map, which itself supports standalone and flexible re-localization. The word “structure” comes from two aspects, as the frame stores the environmental structure information (in the depth form), and the map frames are sparsely structured with special consideration for efficient association and reducing redundancy. Based on the concept, SF-Loc provides a pipeline for the auto-generation of visual structure frames through multi-sensor dense bundle adjustment (MS-DBA), as well as an implementation for coarse-to-fine map-based localization. The contributions of this work are listed as follows:
@@
3) Real-world, cross-season experiments are conducted to validate different phases of the system.
4) The code is made open-source111 https://github.com/GREAT-WHU/SF-Loc to benefit the community.
@@ II-A VSLAM with multi-sensor fusion
Visual SLAM, which has been extensively studied in the past decade, is considered an effective and low-cost solution for relative pose estimation and mapping. Considering the form of visual information, point features[8, 13], line features[14] and direct photometric errors[15] have been exploited. By integrating IMU[16, 17, 18], GNSS[19, 3], and other sensors[20], the practicality of VSLAM can be improved in terms of robustness, continuity, accuracy and geo-referencing functionality. On this basis, multi-session and multi-agent SLAM systems are developed[21, 22]. Learning-based VSLAM frameworks are also investigated in recent implementations, covering neural representations[23, 24], data-driven association[25, 26, 27] and end-to-end pipelines[28, 29].
For map reuse of multi-sensor integrated VSLAM, obtaining globally smoothed pose estimation in the geographic coordinate system is important, as it impacts the absolute accuracy of the map and the practicality of user-side localization.
@@ II-B Visual place recognition
Visual place recognition (VPR) is a long-existing problem, which plays a crucial role in loop closure, multi-session mapping, re-localization and so on. Early approaches use handcrafted models (like bag-of-words[30] and VLAD[31]) to aggregate local feature descriptors, thus to evaluate the scene similarity. To overcome the long-term change challenge, sequential information is utilized[32].
In the last decade, deep neural network (DNN)-based methods, like NetVLAD[33], have greatly improved the practicality and accuracy of VPR. Following this trend, TransVPR[34] uses the transformer to incorporate multi-level attention mechanisms and improve the VPR performance. CosPlaces[35] transforms the place recognition problem from contrastive learning to a classification problem, which enables efficient training on 10-million-level large-scale datasets and is further enhanced with better viewpoint robustness[36]. AnyLoc[37] utilizes the visual foundation model to achieve ubiquitous VPR for different scenarios. It is noted that, most existing methods don’t consider comprehensively the spatial distribution of database images and user-side relative poses, which could serve as powerful information to improve the stability of VPR.
@@ II-C Fine-grained visual re-localization
For high-accuracy visual re-localization (e.g. decimeter- or centimeter-level), it is generally needed to perform feature- or instance-level association between user perception and the map elements. Many of these methods follow a coarse-to-fine pipeline. In ORB-SLAM3[8] and VINS-Mono[16], re-localization is performed through bag-of-words-based place recognition and pointwise matching between new feature observations and the point cloud map. Such pipeline is further improved in HLoc[9], where the power of deep learning is fully utilized in both mapping and localization phases. Besides, monocular depth prediction is exploited in [39] to achieve map-free visual re-localization.
Along with these methods, other approaches try to exploit higher-level information for map matching. Instance-level object modeling is integrated into classic VSLAM systems to overcome viewpoint variation and achieve long-term consistency[40, 41, 42, 43]. In recent works, bird’s-eye view (BEV) features are utilized to associate with priori vector maps to achieve accurate pose estimation, both for outdoor[44] and indoor scenarios[45]. The tradeoff of semantic-based re-localization is the loss of low-level, sometimes subtle but distinguishable, visual cues, which might lead to higher ambiguity.
@@ III System Overview
In the mapping phase, multi-sensor information is used to generate the feature-free map frames with recovered dense depth. Global co-visibility graph optimization is performed periodically in a low frequency to achieve global optimal pose estimation. On this basis, the local co-visibility is checked to keep the map sparsity and to achieve incremental mapping. Besides, a DNN-based VPR model is employed to compute the global descriptors of map frames.
In the localization phase, deep visual place recognition and feature matching models are exploited to achieve coarse-to-fine localization. During this process, multi-frame information and the map frame distribution are fully considered to improve the recall rate and the accuracy.
One principle behind the proposed map representation is to keep the visual information raw and non-encoded. The first advantage of doing so is that it ensures users have flexibility when using the map, without being constrained by any intermediate forms of information. The second advantage is that, the compression of a raw image is effective by applying mature techniques like JPEG[46], without losing the dominant geometric or appearance information.
@@ IV-A Multi-Sensor Dense Bundle Adjustment
The concept of dense bundle adjustment (DBA) is proposed in[26], which is inherited in this work. Specifically, sequential images are fed into a convolutional neural network (CNN) with gated recurrent units (GRUs) to compute dense optical flows among co-visible image pairs. To explain this, consider the rigid flow produced by dense projection of one image pair (source frame $i$ projected to target frame $j$)
@@
where $\Pi_{c}$ is the camera projection model, $\mathbf{T}_{ij}$ is the relative camera pose, $\mathbf{u}_{i}$ denotes the $n=(H/8)\times(W/8)$ grid-like 2D points in the image frame, $\bm{\lambda}_{i}$ is the inverse depth map.
The optical flow module takes the CNN-encoded image feature map and the rigid flow as input and outputs the corrected optical flow (in the residual form) together with the weight $\mathbf{w}_{ij}$, which leads to the following error equation
@@
The residual-form dense optical flows are treated as re-projection errors, which are further used for bundle adjustment to simultaneously estimate the poses and dense depths of the images. Furthermore, the estimated poses and depths are recurrently fed back into the optical flow module, thus to iteratively refine all the geometric information. This process is end-to-end trainable, which shows good generalization in real-world scenes with synthetic training[26].
In our early work[48], we demonstrate the possibility of tightly integrating DBA with multi-sensor information. To be specific, the Hessian form of (2) is derived and stacked for all projections from frame $i$, leading to
@@
where $\bm{\xi}_{i,1,\cdots,N}$ is the stacked Lie algebras of camera poses.
@@
$\mathbf{x}_{k}=\left(\begin{matrix}\mathbf{T}^{w}_{b_{k}}&\mathbf{v}^{w}_{b_{k}}&\mathbf{b}_{k}\end{matrix}\right),k\in\mathcal{K}$ (5)
$\mathbf{T}^{w}_{b_{k}}=\left[\begin{matrix}\mathbf{R}^{w}_{b_{k}}&\mathbf{t}^{w}_{b_{k}}\\ 0&1\end{matrix}\right]\in SE(3),\ \ \mathbf{b}_{k}=\left(\begin{matrix}\mathbf{b}_{a,k}&\mathbf{b}_{g,k}\end{matrix}\right)$ (6)
@@
where $\mathbf{H}_{c}$, $\mathbf{v}_{c}$ are the Hessian matrix and vector of DBA (stacked from $\mathbf{H}_{c,i}$, $\mathbf{v}_{c,i}$) with all the depth states eliminated, $\mathbf{x}_{c}$ is the related camera poses. To associate the camera poses with the IMU-centered pose states, the IMU-camera extriniscs are used to transform (7), as detailed in[48].
@@
$\begin{split}\min\ \ \mathbf{E}_{m}(\mathbf{x})+\mathbf{E}_{c}(\mathbf{x}_{c})&\\ +\sum_{k\in\mathcal{K}}\left\|\mathbf{r}_{b}(\mathbf{x}_{k},\mathbf{x}_{k+1})\right\|^{2}_{\bm{\Omega}_{b}}&+\sum_{k\in\mathcal{G}}\rho_{C}\left(\left\|\mathbf{r}_{g}(\mathbf{x}_{k})\right\|^{2}_{\bm{\Omega}_{g}}\right)\end{split}$ (10)
where $\mathbf{E}_{m}(\mathbf{x})$ is the marginalization term in the quadric cost function form, $\bm{\Omega}_{b}$ and $\bm{\Omega}_{g}$ are uncertainties of the factors, $\rho_{C}(\cdot)$ is the Cauchy loss function.
@@
then the estimated poses and depths are fed back into the optical flow module to refine the optical flows and adjust the weights. For every projection, the optical flow module will be recurrently called for several times until mature. The overall process simultaneously adjust the pose estimation, the dense data association and the depths, trying to achieve consistency of the above information.
An issue that should be considered is the obtainment of global optimal pose estimation. This is important both for monocular visual-inertial odometry (due to the dynamic-dependent observability) and GNSS-integrated navigation (due to intermittent GNSS availability), which can’t be simply achieved in a sliding-window estimator. In our system, we maintain a global factor graph, which collects the marginalized factors from the sliding-window factor graph. The objective of the global factor graph is written as follows
$\begin{split}\min\ \ \sum_{k\in\mathcal{K}_{\text{all}}}\!\!\mathbf{E}_{c,k}(\mathbf{x}_{c,k})\!+\!\!\sum_{k\in\mathcal{K}_{\text{all}}}\!\!\left\|\mathbf{r}_{b}(\mathbf{x}_{k},\mathbf{x}_{k+1})\right\|^{2}_{\bm{\Omega}_{b}}\!\!\\ +\!\!\sum_{k\in\mathcal{G}_{\text{all}}}\!\!\rho_{C}\left(\left\|\mathbf{r}_{g}(\mathbf{x}_{k})\right\|^{2}_{\bm{\Omega}_{g}}\right)\end{split}$ (12)
@@ IV-B Map Construction and Maintenance
After the global pose optimization, we extract the frame pose from the global graph, pack the RGB image (with JPEG compression), the low-resolution depth map and the global descriptor produced by some VPR model together to construct a visual structure frame, then insert it into the map. To make the map lightweight, it is important to keep the sparsity of the visual structure frames without sacrificing too much of the effectiveness of user-side re-localization. Besides, considering the application scenario of multi-session mapping, the system should support incremental update.
@@ V-A Multi-Frame Place Recognition
We use lightweight CNN-based methods (e.g., [33, 52, 35, 36]) to compute global descriptors and to evaluate the similarity between the query image and map frames. On this basis, we develop a model-driven technique to integrate sequential queries and map frame distribution to improve the robustness and stability of the coarse localization.
@@
To avoid the ambiguity of single-shot VPR, we propose an metric called spatially smoothed similarity (SSS) that combines the information of appearance similarity, map distribution and user-side relative poses. To be specific, we set particles around every landmark frames in the map, with every particle to be a 2D pose $(x,y,\theta)$ that represents a candidate of the query pose. With the help of the user-side odometry, we compute the virtual trajectory of every particle, following
$\left(\mathbf{T}^{i}\right)_{4\times 4\times L}=\mathbf{T}^{w}_{p_{i}}\left\{\mathbf{T}^{c_{k}}_{c_{k+1-j}}\right\}_{j=1,2,\cdots,L},{i\in\mathcal{P}}$ (15)
where $\mathcal{P}$ is the particle candidate set, $L$ is the window size used for multi-frame VPR.
@@
$\left({\text{idx}}^{i}\right)_{L\times 1}=\text{db}.\text{nearest}(\mathbf{T}^{i}),\ i\in\mathcal{P}$ (16)
$\left(\bm{\sigma}^{i}\right)_{L\times 1}=\bm{\sigma}\left[\text{idx}^{i}\right],\ i\in\mathcal{P}$ (17)
where db is the map database organized by a spatial KD-Tree, $\bm{\sigma}$ is the $L\times M$ similarity matrix which stores the L2 similarities between the $L$ query frames and all $M$ frames in the map. The similarity computation is performed every time a new query frame comes and stored in the memory.
Then, we compute the SSS distance by averaging the L2 distances
$\sigma^{i}_{\text{SSS}}=||\bm{\sigma}^{i}||/\sqrt{L},\ i\in\mathcal{P}$ (18)
@@
The rationality of SSS is based on the assumption that the historical images of a correctly recalled query should also exhibit high similarity to their spatially adjacent map frames. The geographical distribution of map frames (which is obtained in the mapping phase), along with user-side pose information, is implicitly incorporated to the metric in nearest neighbor search. This technique can be easily adopted to common VPR models without extra layers or finetuning, transforming single-frame VPR similarities to multi-frame ones.
@@
Figure 5: Illustration of the multi-frame place recognition. Particles that indicate the query pose are set around map frames. The user-side trajectory is used for nearest searching to associate with the map frames, then the SSS distance is evaluated, which combines the information of multiple image pairs.